# Condensing Against Online Adversaries

### v1

@@ 6.1 Low-entropy oNOSF source to uniform using two-source-extractors
7 Online influence
Our results from Section 4 provide an almost complete picture on the possibility of condensing from uniform ${\left(g,\ell,n\right)}$-oNOSF source, when the length of each symbol (i.e., $n$) is a large enough constant. In this section, we focus on the class of oNOBF sources , which corresponds to uniform ${\left(g,\ell,n=1\right)}$-oNOSF source and, as noted before, provides the adversary with the most amount of power. We prove lower bounds on the possibility of extraction for oNOBF sources . To this end, we introduce a new, natural notion of influence of Boolean functions, which we call online influence.
We initiate a study of online influence, that we believe is of independent interest. We formally define the notion and discuss some basic properties in Section 7.1. We establish tight bounds on the online influence for general functions, including a Poincaré style inequality, in Section 7.2. We provide an example exhibiting a separation between maximum (standard) influence and online influence in Section 7.3. Finally, we present lower bounds for extraction and condensing from oNOBF sources in Section 7.4.
@@ 7.1 Basic Properties
In this section, for a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we will freely use commas to indicate concatenation in its input. For example, for $x\in\{0,1\}^{i-1}$ and $y\in\{0,1\}^{{\ell}-i+1}$, we write $f(x,1,y)$ to indicate $f$ applied to the tuple $(x_{1},\dots,x_{i-1},1,y_{1},\dots,y_{{\ell}-i+1})$.
@@
It is easy to see that for any $f:\{0,1\}^{\ell}\to\{0,1\}$, and any $i\in{\ell}$, $\operatorname{\mathbf{oI}}_{i}(f)\leq\operatorname{\mathbf{I}}_{i}(f)$. Further, they are the same for the last bit: $\operatorname{\mathbf{I}}_{\ell}[f]=\operatorname{\mathbf{oI}}_{\ell}[f]$.
@@
Using the monotonicity of $f$, note that for any $x\in\{0,1\}^{i-1}$ and any $y\in\{0,1\}^{{\ell}-i}$, $f(x,1,y)\geq f(x,0,y)$. Thus, $\operatorname{\mathbf{oI}}_{i}[f]=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1},y\sim\mathbf{U}_{{\ell}-i}}[f(x,1,y)]-f(x,0,y)]=\operatorname{\mathbf{I}}_{i}(f)$.
@@ 7.2 A Poincaré inequality for online influence
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(f)\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{{\ell}\operatorname*{Var}(f)}$.
Before proving the above result, we observe that MAJORITY and PARITY functions provide tight examples (up to constants) for the upper and lower bound respectively for Theorem 7.5.
@@
The PARITY function on ${\ell}$ bits :$\bigoplus_{\ell}:\{0,1\}^{\ell}\to\{0,1\}$ for $i\in[{\ell}-1]$ has online influence $\operatorname{\mathbf{oI}}_{i}[\bigoplus_{\ell}]=0$, while $\operatorname{\mathbf{oI}}_{\ell}[\bigoplus_{\ell}]=1$. Thus, PARITY meets the lower bound of Theorem 7.5. We note that this is starkly different from regular influence where $\operatorname{\mathbf{I}}_{i}[\bigoplus_{n}]=1$ for all $i$.
To prove Theorem 7.5, we will use Boolean Fourier analysis. Recall that any Boolean function $f:\{0,1\}^{\ell}\rightarrow\{0,1\}$ has a unique Fourier expansion: $f(x)=\sum_{S\subseteq[{\ell}]}\widehat{f}(S)\chi_{S}(x)$, where $\chi_{S}(x)=(-1)^{\sum_{i\in S}x_{i}}$ and $\widehat{f}(S)=\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{\ell}}[(-1)^{f(y)}\chi_{S}(y)]$. Also recall that $\widehat{f}(\emptyset)=\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]$, $\operatorname*{Var}(f)=\sum_{S\subseteq[{\ell}],S\neq\emptyset}\widehat{f}(S)^{2}$, and for any $S\neq T$, $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[\chi_{S}(x)\chi_{T}(x)]=0$. For more background, we refer the reader to the excellent book by O’Donnell [o2014analysis].
@@
We first show the lower bound $\operatorname{\mathbf{oI}}[f]\geq\operatorname*{Var}(f)$. We have,
$\geq\sum_{i=1}^{\ell}\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$ (Lemma 7.7)
$=\sum_{\begin{subarray}{c}S\subseteq[{\ell}]\\ S\neq\varnothing\end{subarray}}\widehat{f}(S)^{2}$
$=\operatorname*{Var}(f).$
@@
We first prove the inequality $\operatorname{\mathbf{oI}}_{i}(f)\geq\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$. Note that $\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)$ is in $[0,1]$ since $\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,b}(y)]$ is in $[0,1]$ for all $x\in\{0,1\}^{i-1},b\in\{0,1\}$. Moreover, for any $x\in\{0,1\}^{i-1}$ we have $\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,1}(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,0}(y)]\right\rvert=\left\lvert\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)\right\rvert$ by 7.8.
@@
$=\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(T)\widehat{f}(S)\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}[\chi_{T\setminus\{i\}}(x)\chi_{S\setminus\{i\}}(x)]$
@@
$=\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$ (derived above)
@@
We begin by defining the restriction $f|_{x,b}(y)=f(x,b,y)$ for $x\in\{0,1\}^{i-1}$, $b\in\{0,1\}$, and $y\in\{0,1\}^{{\ell}-i}$. Thus, can we rewrite $\operatorname{\mathbf{oI}}_{i}[f]$ as
$\operatorname{\mathbf{oI}}_{i}[f]$ $=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,1}(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,0}(y)]\right\rvert\right].$ (1)
@@
This allows us to easily take the expectation of $f|_{x,b}$ and get a result in terms of the Fourier coefficients of $f$.
$\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}\left[f|_{x,b}(y)\right]$ $=\widehat{f|_{x,b}}(\varnothing)$
@@
We now plug this in to our definition of $\operatorname{\mathbf{oI}}_{i}[f]$ in Equation 1 to get a simplified expression.
$\operatorname{\mathbf{oI}}_{i}[f]$ $=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,1}(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[f|_{x,0}(y)]\right\rvert\right]$
$=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\left(-\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)+\sum_{T\subseteq[i-1]}\widehat{f}(T)\chi_{T}(x)\right)-\left(\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{b\}}(x)+\sum_{T\subseteq[i-1]}\widehat{f}(T)\chi_{T}(x)\right)\right\rvert\right]$
@@ 7.3 A tight example for maximum online influence
For a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname{\operatorname{\mathbf{oI}}_{\max}}[f]\geq\operatorname*{Var}(f)/{\ell}$.
@@
By Theorem 7.5 we have that $\operatorname{\mathbf{oI}}[f]=\sum_{i=1}^{\ell}\operatorname{\mathbf{oI}}_{i}[f]\geq\operatorname*{Var}(f)$, and the conclusion follows via an averaging argument. ∎
@@
For $i\in[\log{\ell}]$, no matter what the value of the $i$-th bit of $\operatorname{Addr}_{\ell}$ is set to, the output bit will be a uniform bit, so we immediately get that $\operatorname{\mathbf{oI}}_{i}[f]=0$. Formally, we see this by computing
$=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\frac{1}{2}-\frac{1}{2}\right\rvert\right]$
$=0.$
For $i\in\{\log{\ell}+1,\dots,m\}$, the $i$-th bit only has control if it’s selected by the first $\log{\ell}$ address bits, meaning it has a $1/{\ell}$ chance of controlling the output (and otherwise the output is uniform). Hence, $\operatorname{\mathbf{oI}}_{i}[f]=\frac{1}{{\ell}}$. ∎
@@
Let $\mathsf{Ext}$ be defined as follows: From the first block, use the first $\log(\ell-1)$ bits and interpret them as an index $j\in[\ell-1]$. Then, output block with index $j+1$. For a source $\mathbf{X}$ with first block controlled by an adversary, the output will be truly uniform and for a source $\mathbf{X}$ with adversary controlling one of the last $\ell-1$ blocks, that block will be outputted with probability $\frac{1}{\ell-1}$ while a uniform block will be outputted otherwise. This makes our total error at most $\frac{1}{\ell-1}$ as desired. ∎
@@ 7.4 Online influence of sets and extraction lower bounds
To prove lower bounds on extraction for oNOBF sources , we extend the definition of online influence to set in the natural way.
Definition 7.14 (Online influence of sets and online-resilient functions).
For any function $f:\{0,1\}^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner, starting with sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly, and $\mathbf{A}$ picks the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$; next the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly, and $\mathbf{A}$ sets the value of $x_{i_{2}}$ based on all set variables so for, and so on. Define the advantage of $\mathcal{A}$, $\textrm{adv}_{f,B}(A)=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell}]|$. Then, $I_{B}(f)=\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(A)\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ (that control bits in $B$).
@@
Online-resilient functions can be used to construct extractors for oNOBF sources .
Lemma 7.15.
Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be a $(b,\varepsilon_{1})$-online-resilient function with the property that $\left\lvert f(\mathbf{U}_{\ell})-\mathbf{U}_{1}\right\rvert\leq\varepsilon_{2}$. Then $f$ can extract from ${\left(g=\ell-b,\ell\right)}$-oNOBF sources with error at most $\varepsilon_{1}+\varepsilon_{2}$.
@@
Consider a ${\left(g=\ell-b,\ell\right)}$-oNOBF source $\mathbf{X}$. Recall that $\mathbf{X}$ is created by choosing some set of bad indices $B$ of size $b$, letting the bits in $\overline{B}$ be uniform, and finally setting the bits in $B$ adversarially while only depending on uniform bits to the left of them. Using the triangle inequality for total variation distance, we get that
@@
Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[f(x)=1]=\alpha$. Then for any $1\geq\beta>\alpha$, there exists a coalition $B\subseteq[{\ell}]$ such that $\operatorname{\mathbf{oI}}_{B}(f)\geq\beta-\alpha$, where $\left\lvert B\right\rvert\leq\gamma n$ and $\gamma=\frac{\beta-\alpha}{\alpha(1-\beta)}$.
@@
We greedily collect the bits with the most online influence and add them to $B$ until our goal of $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}|_{\overline{B}}}[f|_{\overline{B}}(x)=1]\geq\beta$ is achieved. Our first step is as follows: let $B_{0}=\varnothing$, $f_{0}=f$, and $i_{1}=\operatorname*{\arg\!\max}_{i\in[{\ell}]}\{\operatorname{\mathbf{oI}}_{i}[f]\}$. Corollary 7.10 tells us that $\operatorname{\mathbf{oI}}_{i_{1}}\geq\operatorname*{Var}(f_{0})/n$. Recall that if $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[f(x)=1]=p$ then $\operatorname*{Var}(f)=p(1-p)$. Because we have not yet achieved our goal of $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}|_{\overline{B}}}[f|_{\overline{B}}(x)=1]\geq\beta$, we have that $\operatorname*{Var}(f_{0})\geq\alpha(1-\beta)$. Thus, we collect $i_{1}$ as $B_{1}=\{i_{1}\}$, let $f_{1}=f_{0}|_{\overline{B_{1}}}$ and see that $\operatorname*{\mathbb{E}}_{x}[f_{1}(x)]\geq\operatorname*{\mathbb{E}}_{x}[f_{0}]+\operatorname{\mathbf{oI}}_{i_{1}}[f_{0}]\geq\alpha+\frac{\alpha(1-\beta)}{{\ell}}$.
@@
$\geq\alpha+\frac{\alpha(1-\beta)}{{\ell}}(t-1)+\frac{\alpha(1-\beta)}{{\ell}}$
$=\alpha+\frac{\alpha(1-\beta)}{{\ell}}\cdot t.$
We repeat this process until $\Pr_{x}[f_{t}(x)=1]\geq\beta$. Therefore, the number of steps is the smallest $b$ such that $\alpha+\frac{\alpha(1-\beta)}{{\ell}}\cdot b\geq\beta$, meaning that the number of steps is at most $b\leq{\ell}\cdot\frac{\beta-\alpha}{\alpha(1-\beta)}$. We let $B=B_{b}$ and get the desired coalition. ∎

### v2

@@ 6.1 Low-Entropy oNOSF Source to Uniform Using Two-Source-Extractors
7 Online Influence and Extraction Lower Bounds
Towards proving lower bounds on the possiblity of extraction from oNOSF sources , we introduce a new, natural notion of influence of Boolean functions, which we call online influence. For simplicity, we first start by considering the class of oNOBF sources , which corresponds to uniform ${\left(g,\ell,n=1\right)}$-oNOSF sources. We formally define the notion for Boolean functions and discuss some basic properties in Section 7.1. We establish tight bounds on the online influence for general functions, including a Poincaré style inequality, in Section 7.2. We provide an example exhibiting a separation between maximum (standard) influence and online influence in Section 7.3. Finally, in Section 7.4, we extend the definition of online influence to subsets of coordinates (and functions from $\Sigma^{n}\rightarrow\{0,1\}^{m}$, for arbitrary alphabet $\Sigma$). This allows us to prove the required lower bounds on extraction (and condensing) from oNOSF sources .
Notation: For convenience, we introduce some notation that we use for the rest of this section. For any bit $b\in\{0,1\}$, let $e(b)=(-1)^{b}$. For any Boolean function $f:\{0,1\}^{\ell}\rightarrow\{0,1\}$, let $e(f)$ denote the function $e(f)(x)=(-1)^{f(x)}$.
@@ 7.1 Basic Properties
In this section, for a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we will freely use commas to indicate concatenation in its input. For example, for $x\in\{0,1\}^{i-1}$ and $y\in\{0,1\}^{{\ell}-i}$, we write $f(x,1,y)$ to indicate $f$ applied to the tuple $(x_{1},\dots,x_{i-1},1,y_{1},\dots,y_{{\ell}-i})$.
@@
It is easy to see that for any $f:\{0,1\}^{\ell}\to\{0,1\}$, and any $i\in{\ell}$, we have $\operatorname{\mathbf{oI}}_{i}(f)\leq\operatorname{\mathbf{I}}_{i}(f)$. Further, they are the same for the last bit: $\operatorname{\mathbf{I}}_{\ell}[f]=\operatorname{\mathbf{oI}}_{\ell}[f]$.
@@
Using the monotonicity of $f$, note that for any $x\in\{0,1\}^{i-1}$ and any $y\in\{0,1\}^{{\ell}-i}$, $f(x,1,y)\geq f(x,0,y)$. Thus, $\operatorname{\mathbf{oI}}_{i}[f]=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1},y\sim\mathbf{U}_{{\ell}-i}}[f(x,1,y)-f(x,1,y)]=\operatorname{\mathbf{I}}_{i}(f)$. ∎
@@ 7.2 A Poincaré Inequality for Online Influence
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(e(f))\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{{\ell}\operatorname*{Var}(e(f))}$.
Before proving the above result, we observe that the MAJORITY and PARITY functions provide tight examples for the upper and lower bound respectively for Theorem 7.5.
@@
The PARITY function on ${\ell}$ bits :$\bigoplus_{\ell}:\{0,1\}^{\ell}\to\{0,1\}$ for $i\in[{\ell}-1]$ has online influence $\operatorname{\mathbf{oI}}_{i}[\bigoplus_{\ell}]=0$, while $\operatorname{\mathbf{oI}}_{\ell}[\bigoplus_{\ell}]=1$. Thus, PARITY meets the lower bound of Theorem 7.5. We note that this is starkly different from regular influence where $\operatorname{\mathbf{I}}_{i}[\bigoplus_{\ell}]=1$ for all $i$.
To prove Theorem 7.5, we will use Boolean Fourier analysis. For any $f:\{0,1\}^{n}\to\{0,1\}$, $e(f)$ has a unique Fourier expansion given by: $e(f(x))=\sum_{S\subseteq[{\ell}]}\widehat{f}(S)\chi_{S}(x)$, where $\chi_{S}(x)=(-1)^{\sum_{i\in S}x_{i}}$ and $\widehat{f}(S)=\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{\ell}}[e(f)(y)\chi_{S}(y)]$. [^8] For simplicity of notation, we use $\widehat{f}(S)$ for $\widehat{e(f)}(S)$. Also recall that $\widehat{f}(\emptyset)=\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{n}}[e(f)(x)]$, $\operatorname*{Var}(e(f))=\sum_{S\subseteq[{\ell}],S\neq\emptyset}\widehat{f}(S)^{2}$, and for any $S\neq T$, $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[\chi_{S}(x)\chi_{T}(x)]=0$. For more background, we refer the reader to the excellent book by O’Donnell [o2014analysis].
@@
We start with the lower bound. We have,
$\operatorname{\mathbf{oI}}[f]=\sum_{i=1}^{\ell}\operatorname{\mathbf{oI}}_{i}[f]\geq\sum_{i=1}^{\ell}\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}=\sum_{\begin{subarray}{c}S\subseteq[{\ell}]\\ S\neq\varnothing\end{subarray}}\widehat{f}(S)^{2}=\operatorname*{Var}(e(f)),$
where the inequality uses Lemma 7.7.
@@
We first prove the inequality $\operatorname{\mathbf{oI}}_{i}(f)\geq\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(S)^{2}$. Since for any $x\in\{0,1\}^{i-1}$ we have $\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,1})(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,0})(y)]\right\rvert=2\left\lvert\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)\right\rvert$ by 7.8, and the fact that $\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,b})(y)]$ is in $[-1,1]$ for all $x\in\{0,1\}^{i-1},b\in\{0,1\}$, it follows that $\left|\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)\right|$ is in $[0,1]$.
@@
$=\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\sum_{\begin{subarray}{c}S\subseteq[i]\\ S\ni i\end{subarray}}\widehat{f}(T)\widehat{f}(S)\cdot\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}[\chi_{T\setminus\{i\}}(x)\chi_{S\setminus\{i\}}(x)]$
@@
[…]
@@
We begin by defining the restriction $f|_{x,b}(y)=f(x,b,y)$ for $x\in\{0,1\}^{i-1}$, $b\in\{0,1\}$, and $y\in\{0,1\}^{{\ell}-i}$. Thus, we can rewrite $\operatorname{\mathbf{oI}}_{i}[f]$ as
$\operatorname{\mathbf{oI}}_{i}[f]$ $=\frac{1}{2}\cdot\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,1})(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,0})(y)]\right\rvert\right].$ (1)
@@
Thus, we have
$\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}\left[e(f|_{x,b})(y)\right]$ $=\widehat{f|_{x,b}}(\varnothing)$
@@
We now plug this in to our definition of $\operatorname{\mathbf{oI}}_{i}[f]$ in Equation 1 to get a simplified expression. Recalling the fact that for any $x\in\{0,1\}^{n}$, $f(x)=(1-e(f)(x))/2$, we have
$\operatorname{\mathbf{oI}}_{i}[f]$ $=\frac{1}{2}\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,1})(y)]-\operatorname*{\mathbb{E}}_{y\sim\mathbf{U}_{{\ell}-i}}[e(f|_{x,0})(y)]\right\rvert\right]$
$=\frac{1}{2}\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{i-1}}\left[\left\lvert\left(-\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{i\}}(x)+\sum_{T\subseteq[i-1]}\widehat{f}(T)\chi_{T}(x)\right)-\left(\sum_{\begin{subarray}{c}T\subseteq[i]\\ T\ni i\end{subarray}}\widehat{f}(T)\chi_{T\setminus\{b\}}(x)+\sum_{T\subseteq[i-1]}\widehat{f}(T)\chi_{T}(x)\right)\right\rvert\right]$
@@ 7.3 A Tight Example for Maximum Online Influence
For a function $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname{\operatorname{\mathbf{oI}}_{\max}}[f]\geq\operatorname*{Var}(e(f))/{\ell}$.
@@
By Theorem 7.5 we have that $\operatorname{\mathbf{oI}}[f]=\sum_{i=1}^{\ell}\operatorname{\mathbf{oI}}_{i}[f]\geq\operatorname*{Var}(e(f))$, and the conclusion follows via an averaging argument. ∎
@@
For $i\in[\log{\ell}]$, no matter what the value of the $i$-th bit of $\operatorname{Addr}_{\ell}$ is set to, the output bit will be a uniform bit, so we immediately get that $\operatorname{\mathbf{oI}}_{i}[f]=0$. For $i\in\{\log{\ell}+1,\dots,m\}$, the $i$-th bit only has control if it’s selected by the first $\log{\ell}$ address bits, meaning it has a $1/{\ell}$ chance of controlling the output (and otherwise the output is uniform). Hence, $\operatorname{\mathbf{oI}}_{i}[f]=\frac{1}{{\ell}}$. ∎
@@
Let $\mathsf{Ext}$ be defined as follows: From the first block, use the first $\log(\ell-1)$ bits and interpret them as an index $j\in[\ell-1]$. Then, output the block with index $j+1$. For a source $\mathbf{X}$ with first block controlled by an adversary, the output will be truly uniform and for a source $\mathbf{X}$ with adversary controlling one of the last $\ell-1$ blocks, that block will be outputted with probability $\frac{1}{\ell-1}$ while a uniform block will be outputted otherwise. This makes our total error at most $\frac{1}{\ell-1}$ as desired. ∎
@@ 7.4 Online Influence of Sets and Extraction Lower Bounds
For convenience we restate the definition of online influence of sets of coordinates.
Definition 7.14 (Online influence, Definition 1.7 restated).
For any function $f:\Sigma^{\ell}\to\{0,1\}$, and any $B\subset[{\ell}]$, where $B=\{i_{1}<i_{2}<\ldots<i_{k}\}$, define $\operatorname{\mathbf{oI}}_{B}(f)$ as follows: an online adversary $\mathcal{A}$ samples a distribution $\mathbf{X}$ in online manner. It starts by sampling the variables $x_{1},x_{2},\ldots,x_{i_{1}-1}$ independently and uniformly from $\Sigma$, then picking the value of $x_{i_{1}}$ depending on $x_{<i_{1}}$. Next, the variables $x_{i_{1}+1},\ldots,x_{i_{2}-1}$ are sampled independently and uniformly from $\Sigma$, and $\mathcal{A}$ sets the value of $x_{i_{2}}$ based on all set variables so for, and so on. Define the advantage of $\mathcal{A}$ to be $\textrm{adv}_{f,B}(\mathcal{A})=|\operatorname*{\mathbb{E}}[f(\mathbf{X})]-\operatorname*{\mathbb{E}}[f(\mathbf{U}_{\ell})]|$. Then $\operatorname{\mathbf{oI}}_{B}(f)$ is defined to be $\max_{\mathcal{A}}\{\textrm{adv}_{f,B}(\mathcal{A})\}$, where the maximum is taken over all online adversaries $\mathcal{A}$ that control the bits in $B$.
@@
In the special case where $\Sigma=\{0,1\}$ and we are considering the online influence of a single coordinate, the definition simplifies nicely.
Definition 7.15.
and the total online influence is
Online-resilient functions are equivalent to extractors (with $1$ output bit) for oNOSF sources .
Lemma 7.16 (online-resilient functions yield extractors).
Let $f:\Sigma^{\ell}\to\{0,1\}$ be a $(b,\varepsilon_{1})$-online-resilient function with the property that $\left\lvert f(\mathbf{U}_{\ell})-\mathbf{U}_{1}\right\rvert\leq\varepsilon_{2}$. Then $f$ can extract from ${\left(g=\ell-b,\ell\right)}$-oNOSF sources with error at most $\varepsilon_{1}+\varepsilon_{2}$.
@@
Consider a ${\left(g=\ell-b,\ell\right)}$-oNOSF source $\mathbf{X}$. Recall that $\mathbf{X}$ is created by choosing some set of bad indices $B$ of size $b$, letting the symbols in $\overline{B}$ be uniform, and finally setting the symbols in $B$ adversarially while only depending on uniform symbols to the left of them. Using the triangle inequality for total variation distance, we get that
@@
We note that the other direction is immediate from definitions. If $\mathsf{Ext}:\Sigma^{\ell}\to\{0,1\}$ is an extractor with error $\varepsilon$ for $(g=\ell-b,\ell)$-oNOSF sources , then $\mathsf{Ext}$ is a $(b,2\varepsilon)$-online-resilient function.
Remark 7.18.
@@
Let $f:\{0,1\}^{\ell}\to\{0,1\}$ be such that $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[f(x)=1]=\alpha$. Then for any $1\geq\beta>\alpha$, there exists a coalition $B\subseteq[{\ell}]$ such that $\operatorname{\mathbf{oI}}_{B}(f)\geq\beta-\alpha$, where $\left\lvert B\right\rvert\leq\gamma\ell$ and $\gamma=\frac{\beta-\alpha}{4\alpha(1-\beta)}$.
@@
We greedily collect the bits with the most online influence and add them to $B$ until our goal of $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}|_{\overline{B}}}[f|_{\overline{B}}(x)=1]\geq\beta$ is achieved. Our first step is as follows: let $B_{0}=\varnothing$, $f_{0}=f$, and $i_{1}=\operatorname*{\arg\!\max}_{i\in[{\ell}]}\{\operatorname{\mathbf{oI}}_{i}[f]\}$. Corollary 7.10 tells us that $\operatorname{\mathbf{oI}}_{i_{1}}\geq\operatorname*{Var}(e(f_{0}))/\ell$. Recall that if $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}}[f(x)=1]=p$ then $\operatorname*{Var}(e(f))=4p(1-p)$. Because we have not yet achieved our goal of $\operatorname*{\mathbb{E}}_{x\sim\mathbf{U}_{\ell}|_{\overline{B}}}[f|_{\overline{B}}(x)=1]\geq\beta$, we have that $\operatorname*{Var}(f_{0})\geq 4\alpha(1-\beta)$. Thus, we collect $i_{1}$ as $B_{1}=\{i_{1}\}$, let $f_{1}=f_{0}|_{\overline{B_{1}}}$ and see that $\operatorname*{\mathbb{E}}_{x}[f_{1}(x)]\geq\operatorname*{\mathbb{E}}_{x}[f_{0}(x)]+\operatorname{\mathbf{oI}}_{i_{1}}[f_{0}]\geq\alpha+\frac{4\alpha(1-\beta)}{{\ell}}$.
@@
$\geq\alpha+\frac{4\alpha(1-\beta)}{{\ell}}(t-1)+\frac{4\alpha(1-\beta)}{{\ell}}$
$=\alpha+\frac{4\alpha(1-\beta)}{{\ell}}\cdot t.$
We repeat this process until $\Pr_{x}[f_{t}(x)=1]\geq\beta$. Therefore, the number of steps is the smallest $b$ such that $\alpha+\frac{4\alpha(1-\beta)}{{\ell}}\cdot b\geq\beta$, meaning that the number of steps is at most $b\leq{\ell}\cdot\frac{\beta-\alpha}{4\alpha(1-\beta)}$. We let $B=B_{b}$ and get the desired coalition. ∎